I am in desperate...I can do these problems until I get to the answer and I miss every time, so could someone please give me step by step.
Find an equation of the line containing the given pair of points
(-2, -3) and (-8, -4)


THANKS!

Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(7384)

First, find the slope from:
m = (y2-y1)/(x2-x1)
Given:(-2, -3) and (-8, -4)
.
m = (-4-(-3))/(-8-(-2))
m = (-4+3)/(-8+2)
m = (-1)/(-6)
m = -1/6
.
Now, using the slope (-1/6) above and one of the two given points say (-2, -3) plug it into the point slop form of:
(y-y1) = m(x-x1)
(y-(-3)) = (-1/6)(x-(-2))
y+3 = (-1/6)(x+2)
y+3 = (-1/6)(x) + (-1/6)2
y+3 = (-1/6)(x) + (-1/3)
y = (-1/6)(x) + (-1/3) - 3
y = (-1/6)(x) + (-1/3) - 9/3
y = (-1/6)(x) + (-10/3)
y = (-1/6)(x) - 10/3 (this is your solution in "slope-intercept" form)

Answer by stanbon(75887)

Find an equation of the line containing the given pair of points
(-2, -3) and (-8, -4)
-----
slope : (-4--3) / (-8--2) = -1/-6 = 1/6
---------------
Intercept:
Since the form is y = mx + b and you know y,x, and m, solve for "b".
-3 = (1/6)*-2 + b
-3 = -1/3 + b
b = -9/3 + 1/3 = -8/3
---------------------------
Equation:
y = (1/6)x - (8/3)
=======================
Cheers,
Stan H.